Please help me:
1/3x-9= 2(x-5)

Answer by rothauserc(4718)

1/3x-9= 2(x-5)
multiply both sides of the = by 3
x-27 = 6(x-5)
expand right side of the =
x-27 = 6x-30
add 27 to both sides of the =
x = 6x -3
subtract 6x from both sides of =
-5x = -3
x = 3/5